If cot θ = 2, find the value of all T-ratios of θ.
find

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given function: cot θ= 2


We know that cot θ = base/perpendicular
= 2/1
where k is a positive

By Pythagoras Theorem:

AB^2 = BC^2 + AC^2
= k^2 + 4k^2
= 5k^2
AB = k√5

Find other T-rations using their definitions:
Sin𝜃 = k/(k√5) = 1/√5

cos𝜃 = (2k)/(k√5) = 2/√5

tan θ = sinθ /cosθ = k/(2k) = 1/2

cosec 𝜃= 1/sin𝜃= √5

sec𝜃= 1/cos𝜃= √5/2

Thanks